OpenStudy (bekkah323):

factor 3a^2-12ab-125b^2

OpenStudy (bekkah323):

the book gave me the answer 3(a-9b)(a+5b) i just don't know how they got there

OpenStudy (agent0smith):

Hmm... the book's answer doesn't match the question, it would give 135b^2.

OpenStudy (bekkah323):

your right its not 125 its 135

OpenStudy (bekkah323):

sorry

OpenStudy (agent0smith):

No prob. First notice you can factor out a 3 out of all the terms:\[\Large 3a^2-12ab-135b^2 = 3 (a^2-4ab -45b^2)\]

OpenStudy (bekkah323):

ok

OpenStudy (agent0smith):

Now try to factor whats inside the brackets. Don't worry so much about the a and b, just find two numbers that multiply to -45, and add up to -4.

OpenStudy (agent0smith):

You can always try listing the factors of 45 to see what will work 45 is 3*15, 5*9... -9 and +5 add up to -4. Now you can factor it. You'll just have an a and a b in each one... since -9b*5b gives -45b^2 \[\Large 3 (a^2-4ab -45b^2) = 3 (a - 9b) (a+5b)\]

OpenStudy (bekkah323):

ah ok thank you, that makes sense.

OpenStudy (agent0smith):

Good :)